OpenStudy (anonymous):

how do u do this problem x/5+6=-2

OpenStudy (anonymous):

Subtract 6 from both sides \[\frac{x}{5}+6-6=-2-6\] \[\frac{x}{5}=-8\] Multiply each side by 5 \[\frac{x}{5}*5=-8*5\] \[x=-40\]

OpenStudy (anonymous):

x/5+6=-2

OpenStudy (anonymous):

h/4-8=-5

OpenStudy (anonymous):

h/4-8=-5